OpenStudy (anonymous):

if f(x)=x^2+1and9(x)=3x+1evaluate f(a+1)

OpenStudy (mathmale):

Welcome to OpenStudy! We need some clarification here, along with some tips on presentation of your math expressions. Please leave space (at least one space) between each new word or equation: f(x)=x^2+1and9(x)=3x+1 => f(x)=x^2+1 and 9(x)=3x+1. What do you mean by 9(x)? Please clarify that.